# Towards Lightweight Speaker Verification via Adaptive Neural Network Quantization

### v1

@@ III-A Preliminaries
Network quantization, a common approach for model compression, primarily involves two essential operations: quantization and dequantization. There are two different types of strategies: post-training quantization (PTQ) and quantization-aware training (QAT). PTQ directly applies quantization to pre-trained weights without requiring any further training. Although it is simple and efficient, the performance of quantized models is not satisfactory. On the contrary, QAT introduces a fake node [47] into the computational graph to simulate the quantization process. Through gradient approximation, system can be trained in an end-to-end manner, thereby yielding better performance. Our work is based on the framework of QAT.
@@
where $\alpha$ denotes a learnable scaling factor with full precision.
For a network, we can construct a set of quantization levels $q$ and its associated scaling factor $\alpha$ for every layer within it.
quantization-aware training: It is a technique that takes the effects of quantization on a model’s performance into consideration. During the training process, a full-precision copy of the model weights is retained to mitigate quantization errors. In addition, straight-through estimation (STE) [50] is adopted to directly assign incoming gradients as outgoing gradients for quantization operation (Eq. 5). Once the model is trained, only the quantized weights are utilized for inference.
@@ III-B Adaptive Uniform Precision Quantization
step 3: We empirically observe that the majority of quantization centroid $q$ from step 2 fall within the magnitude range of $10^{-3}\sim 10^{-2}$, potentially causing numerical instability in scaling factor $\alpha$ learning during quantization-aware training. As Fig. 1 displays, we propose two different affine transformations, namely asymmetric and symmetric rescaling, to project quantization centroids from step 2 into $[-1,1]$ for training stability. In asymmetric rescaling, we perform a direct mapping of $q_{min}$ to $-1$ and $q_{max}$ to $1$. The formula is as follows:
$q=\frac{q-q_{min}}{q_{max}-q_{min}}\times 2-1$ (6)
@@
Through the above three steps, our resulting quantization levels $q$ can better align with the weight distribution across all network layers (Fig. 2(d)), thereby enjoying superior representation capability. Combined with quantization-aware training technique, our approach is named as K-Means based Quantization Aware Training (KMQAT).
@@ IV Mixed Precision Quantization
Estimate Hessian Sensitivity: As illustrated in Fig. 3, we initially compute the average trace of Hessian matrix for the pre-trained weights, employing this as the sensitivity metric for each layer. Subsequently, all network layers are re-arranged in a descending according to their average traces. Layers with a large average trace will be given a higher bit width, and conversely for those with a low average trace. This is reasonable because when layers exhibit a significant average trace, indicating a higher vulnerability to quantization impacts, it is becomes to allocate a greater bit precision to mitigate performance degradation.
@@
where $H_{i}$ represents Hessian matrix of the $i$-th layer and $L$ is the total number of network layers. $\overline{\operatorname{Tr}}\left(H_{i}\right)$ denotes the average trace of $H_{i}$. $W_{i}$ is pre-trained weights in the $i$-th layer and $Q_{i}$ is the corresponding quantized weights. $\lVert W_{i}-Q_{i}\rVert_{2}^{2}$ is perturbation incurred by quantization on pre-trained weights.
@@
step 3: After obtaining mixed-precision bit combination $s$, we will perform quantization aware fine-tuning in this step. In uniform precision quantization, the entire network undergoes a single round of quantization and fine-tuning. However, this approach might result in a sub-optimal solution for mixed precision quantization, as the bit width differs among various network layers. Instead, we propose a Multi-Stage Fine-Tuning (MSFT) strategy where only parts of network are quantized and fine-tuned at each stage. Specifically, MSFT progressively quantizes layers based on their bit width. As Fig. 3 shows, it starts with low bits and gradually progresses to high ones. Before going to the next stage, the partially quantized models will be fully trained until convergence. Algorithm 1 provides a detailed description of MSFT.
@@ V-A Static Binary Quantization
where $clip$ function clamps weights between -1 and 1. $round$ is used to map values to the closest integer.
@@ V-B Adaptive Binary Quantization
Unlike static quantization that utilizes a fixed binary integer set $\left\{-1,+1\right\}$, we introduce two adaptive parameters, $\alpha$ and $\beta$, to dynamically generate different binary sets for each layer, as Figure 4 (right) displays. The binarized weight can be derived using the following equation:
$Q=\begin{cases}b_{1}=\beta-\alpha,&w<\beta\\ b_{2}=\beta+\alpha,&w>\beta\end{cases}$ (13)
where $\beta$ is defined as the center of binarized weights. $\alpha$ signifies the deviation from this center. In this case, the binary set contains $\left\{b_{1},b_{2}\right\}$, which varies across different layers.
Next, we will discuss how to determine $\alpha$ and $\beta$ for the $i$-th layer. Starting with weight matrix $W_{i}$, we initially align the center of binary value $\beta$ with the mean of $W_{i}$. As a result, $\beta$ can be obtained via:
$\beta=\frac{1}{d\times k\times k}\sum_{j=0}^{d-1}\sum_{m=0}^{k-1}\sum_{n=0}^{k-1}W_{i}^{j,m,n}$ (14)
where $d$ and $k$ correspond to channel number and kernel size respectively. $j$, $m$ and $n$ are indices for channel $d$, the first and second dimension of kernel size $k$ respectively.
When estimating $\alpha$, Kullback-Leibler divergence (KLD) is utilized to evaluate the similarity between the distributions of binarized and real-valued weights.
@@
We assume that binarized weights adhere to a uniform distribution, which means $P_{b}(b_{1})=P_{b}(b_{2})=1/2$. For real-valued weights, it is typically considered that they follow a Gaussian distribution. To minimize the KL distance in Eq. 15, our empirical observations suggest that $\alpha$ should be at the standard deviation of $W_{i}$, which can be calculated as follows:
$\alpha=\frac{\lVert W_{i}-\beta\rVert_{2}}{\sqrt{d\times k\times k}}$ (16)
During quantization-aware training, our adaptive quantization method enables dynamic update of $\alpha$ and $\beta$ along with real-valued weights within each layer.
@@ VII-A Evaluation on Adaptive Uniform Precision Quantization
The results presented in Table III and IV demonstrate that KMQAT effectively realizes lossless 4-bit uniform precision quantization on both ResNets and DF-ResNets, offering an impressive compression ratio of approximately $\sim$8. For example, when applying KMQAT to ResNet34 and DF-ResNet110, the resulting 4-bit quantized models attain nearly identical performance as opposed to their full-precision versions. This confirms the effectiveness of using k-means clustering for the generation of quantization centroids. Similarly, the performance loss caused by 4-bit compression in larger and deeper networks, such as ResNet101 and DF-ResNet179, is negligible. This highlights the strong generality of our proposed KMQAT method, suggesting its broad applicability across different network architectures. Compared to fixed uniform method, our adaptive one achieves much better performance across various bit precisions, as Fig. 6 shows.
@@
VII-B Evaluation on Mixed Precision Quantization
From Table III and IV, it can be clearly observed that the proposed mixed precision approach significantly outperforms uniform precision quantization, particularly in low-bit scenarios, while maintaining a similar model size and compression ratio. Specifically, bit combinations discovered for ResNet34 and DF-ResNet110 achieve the same level of 3-bit quantization as uniform precision with a slightly smaller model size. When integrated with MSFT, these mixed precision quantized models exhibit superior performance. Similarly, applying 2-bit mixed precision quantization to ResNets and DF-ResNets obtains average performance improvements of 6.2% and 6.5%, respectively. For cases below 2-bit, we explore three additional configurations: 1.7-bit, 1.3-bit and 1.1-bit. In these contexts, mixed precision quantization demonstrates a larger advantage compared to KMQAT. For instance, 1.1-bit equivalent quantization surpasses the performance of 1-bit KMQAT by an average of 14%. In addition, Fig. 6 presents a schematic comparison between uniform and mixed precision quantization, illustrating that the latter achieves a better trade-off on performance and model size for regime below 3-bit. The above findings suggest that the searched bit combination based on Hessian sensitivity analysis is more effective and reasonable than uniform one, thereby boosting the performance of low-bit quantized models. Moreover, MSFT proves to be a powerful strategy for incrementally quanitzing and fine-tuning network in the context of mixed precision quantization.
@@ VII-D3 Binary Quantization
Finally, we analyze the binarized weight histograms of both static and adaptive quantization schemes. The distributions of pre-trained weights, as shown in Fig. 9, display a notable disparity between the first and last convolutional layers. Using a fixed integer set, static 1-bit quantization produces almost the same binarized results in the first and last layers ($0.07$ vs. $0.06$), which fails to accurately match the pre-trained weight distribution. For example, the majority of weights in final layer fall between $[-0.04,0.04]$. However, static quantization yields two binarized weights at ±$\pm 0.06$. In contrast, adaptive quantization can determine binary values for different layers according to their distributions. In the broader first layer, weights are mapped to ±$\pm 0.05$. Meanwhile, for the last layer with a more compact shape, it generates binarized weights of $-0.02$ and $0.02$. This illustrates the superiority of adaptive scheme over the static method.
@@ VIII Conclusion
In this paper, we explore adaptive neural network quantization for lightweight speaker verification. Firstly, we introduce an innovative adaptive uniform precision quantization technique that leverages k-means clustering to dynamically generate quantization centroids specific to each network layer. To improve the performance of low-bit quantized models, a new algorithm that merges mixed precision quantization with a multi-stage fine-tuning (MSFT) strategy is further developed. Finally, we propose two distinct binary quantization schemes tailored for 1-bit scenario: the static and adaptive quantizers. Experimental results demonstrate that both ResNets and DF-ResNets effectively attain 4-bit uniform precision quantization with negligible performance loss. Furthermore, compared to the uniform precision approach, mixed precision quantization not only achieves better performance with a similar model size but also enables the generation of bit combination for any desirable model size. In addition, our newly designed 1-bit quantization schemes significantly enhance the performance of binarized models. The visualization of quantized weight distributions validates the superiority of our proposed quantization methods. Finally, an in-depth comparison with existing lightweight SV systems indicates that our resulting models significantly outperform earlier systems across various model size ranges.

### v3

@@ III-A Preliminaries
Network quantization, a common approach for model compression, primarily involves two essential operations: quantization and dequantization.
@@
where $\alpha$ denotes a learnable scaling factor with full precision. $Q$ is the quantized version of a given full-precision number. For a network, we can construct a set of quantization levels $q$ and its associated scaling factor $\alpha$ for every layer within it.
There exist two different quantization frameworks: post-training quantization (PTQ) and quantization-aware training (QAT). Both frameworks include the quantize and dequantize operations.
post-training quantization: PTQ [49, 50] directly perform quantization to pre-trained weights without requiring any further training. Despite the simplicity, it can cause significant performance degradation.
quantization-aware training: On the contrary, QAT [51, 52, 53] introduces a fake node into the computational graph to simulate the quantization process. Through gradient approximation, system can be trained in an end-to-end manner, thereby recovering accuracy loss caused by quantization. Our work is based on the framework of QAT. Specifically, it is a technique that takes the effects of quantization on a model’s performance into consideration. During the training process, a full-precision copy of the model weights is retained to mitigate quantization errors. In addition, straight-through estimation (STE) [54] is adopted to directly assign incoming gradients as outgoing gradients for quantization operation (Eq. 5). Once the model is trained, only the quantized weights are utilized for inference.
@@
where $\mathcal{L}$ means the loss function. $W$ represents full-precision weights of a network.
@@ III-B Adaptive Uniform Precision Quantization
step 3: We empirically observe that the majority of quantization centroid $q$ from step 2 fall within the magnitude range of $10^{-3}\sim 10^{-2}$, potentially causing numerical instability in scaling factor $\alpha$ learning during quantization-aware training. As Fig. 1 displays, we propose two different affine transformations, namely asymmetric and symmetric rescaling, to project quantization centroids from step 2 into $[-1,1]$ for training stability. Notably, this is the first work to apply an affine transformation to the quantization centroids.
In asymmetric rescaling, we perform a direct mapping of $q_{min}$ to $-1$ and $q_{max}$ to $1$. The formula is as follows:
$q_{rescale}=\frac{q-q_{min}}{q_{max}-q_{min}}\times 2-1$ (6)
@@
Through the above three steps, our resulting quantization levels $q_{rescale}$ can better align with the weight distribution across all network layers (Fig. 2(d)), thereby enjoying superior representation capability. Combined with quantization-aware training technique, our approach is named as K-Means based Quantization Aware Training (KMQAT).
@@ IV Mixed Precision Quantization
Estimate Hessian Sensitivity: As illustrated in Fig. 3, we initially compute the average trace of Hessian matrix for the pre-trained weights, employing this as the sensitivity metric for each layer. Subsequently, all network layers are re-arranged in a descending according to their average traces. Layers with a large average trace will be given a higher bit width, and conversely for those with a low average trace. This is reasonable because when layers exhibit a significant average trace, indicating a higher vulnerability to quantization impacts, it becomes necessary to allocate a greater bit precision to mitigate performance degradation.
@@
where $\Omega_{i}$ signifies the estimated sensitivity of the $i$-th layer, and $\Omega_{s}$ indicates the total sensitivity of a network with $L$ layers. $H_{i}$ represents Hessian matrix of the $i$-th layer. $\overline{\operatorname{Tr}}\left(H_{i}\right)$ denotes the average trace of $H_{i}$. $W_{i}$ is pre-trained weights in the $i$-th layer and $Q_{i}$ is the corresponding quantized weights. $\lVert W_{i}-Q_{i}\rVert_{2}^{2}$ is the second-order perturbation incurred by quantization on pre-trained weights.
@@
step 3: After obtaining mixed-precision bit combination $s$, we will perform quantization aware fine-tuning in this step. In uniform precision quantization, the entire network undergoes a single round of quantization and fine-tuning. However, this approach might result in a sub-optimal solution for mixed precision quantization. Previous studies [58, 59, 55] have revealed that quantization errors can vary significantly across different network layers. Layers with a higher bit width typically exhibit smaller quantization errors, while those with a lower bit width can have much larger errors. Quantizing weights all at once in a single iteration can lead to inappropriate gradient direction during training, causing the model to converge to a sub-optimal local minimum. Instead, we propose a Multi-Stage Fine-Tuning (MSFT) strategy where only parts of network are quantized and fine-tuned at each stage. Specifically, MSFT progressively quantizes layers based on their bit widths. As Fig. 3 shows, it starts with low bits and gradually progresses to high ones. Before going to the next stage, the partially quantized models will be fully trained until convergence. Algorithm 1 provides a detailed description of MSFT.
@@ V-A Static Binary Quantization
where $w^{\prime}$ is an element of the matrix $W_{i}^{\prime}$. $clip$ function clamps weights between -1 and 1. $round$ is used to map values to the closest integer.
@@ V-B Adaptive Binary Quantization
Unlike static quantization that utilizes a fixed binary integer set $\left\{-1,+1\right\}$, we introduce two adaptive parameters, $d$ and $\beta$, to dynamically generate different binary sets for each layer, as Figure 4 (right) displays. The binarized weight can be derived using the following equation:
$Q=\begin{cases}b_{1}=\beta-d,&w<\beta\\ b_{2}=\beta+d,&w>\beta\end{cases}$ (13)
where $\beta$ is defined as the center of binarized weights. $d$ signifies the deviation from this center. In this case, the binary set contains $\left\{b_{1},b_{2}\right\}$, which varies across different layers.
Next, we will discuss how to determine $d$ and $\beta$ for the $i$-th layer. Starting with weight matrix $W_{i}$, we initially align the center of binary value $\beta$ with the mean of $W_{i}$. As a result, $\beta$ can be obtained via:
$\beta=\frac{1}{h\times k\times k}\sum_{j=0}^{h-1}\sum_{m=0}^{k-1}\sum_{n=0}^{k-1}W_{i}^{j,m,n}$ (14)
where $h$ and $k$ correspond to channel number and kernel size respectively. $j$, $m$ and $n$ are indices for channel $h$, the first and second dimension of kernel size $k$ respectively.
When estimating $d$, Kullback-Leibler divergence (KLD) is utilized to evaluate the similarity between the distributions of binarized and real-valued weights.
@@
We assume that binarized weights adhere to a uniform distribution, which means $P_{b}(b_{1})=P_{b}(b_{2})=1/2$. For real-valued weights, it is typically considered that they follow a Gaussian distribution. To minimize the KL distance in Eq. 15, our empirical observations suggest that $d$ should be at the standard deviation of $W_{i}$, which can be calculated as follows:
$d=\frac{\lVert W_{i}-\beta\rVert_{2}}{\sqrt{h\times k\times k}}$ (16)
During quantization-aware training, our adaptive quantization method enables dynamic update of $d$ and $\beta$ along with real-valued weights within each layer.
@@ VII-A Evaluation on Adaptive Uniform Precision Quantization
The results presented in Table III and IV demonstrate that KMQAT effectively realizes lossless 4-bit uniform precision quantization on both ResNets and DF-ResNets, offering an impressive compression ratio of approximately $\sim$8 and $\sim$7 respectively. For example, when applying KMQAT to ResNet34 and DF-ResNet110, the resulting 4-bit quantized models attain nearly identical performance as opposed to their full-precision versions. This confirms the effectiveness of using k-means clustering for the generation of quantization centroids. Similarly, the performance loss caused by 4-bit compression in larger and deeper networks, such as ResNet101 and DF-ResNet179, is negligible. This highlights the strong generality of our proposed KMQAT method, suggesting its broad applicability across different network architectures. Compared to fixed uniform method, our adaptive one achieves much better performance across various bit precisions, as Fig. 6 shows.
@@
VII-B Evaluation on Mixed Precision Quantization
From Table III and IV, it can be clearly observed that the proposed mixed precision approach significantly outperforms uniform precision quantization, particularly in low-bit scenarios, while maintaining a similar model size and compression ratio. Specifically, bit combinations discovered for ResNet34 and DF-ResNet110 achieve the same level of 3-bit quantization as uniform precision with a slightly smaller model size. When integrated with MSFT, these mixed precision quantized models exhibit superior performance. Similarly, applying 2-bit mixed precision quantization to ResNets and DF-ResNets obtains average performance improvements of 6.2% and 6.5%, respectively. For cases below 2-bit, we explore three additional configurations: 1.7-bit, 1.3-bit and 1.1-bit. In these contexts, mixed precision quantization demonstrates a larger advantage compared to KMQAT. For instance, 1.1-bit equivalent quantization surpasses the performance of 1-bit KMQAT by an average of 14% with only a slight increase in model size. In addition, Fig. 6 presents a schematic comparison between uniform and mixed precision quantization, illustrating that the latter achieves a better trade-off on performance and model size for regime below 3-bit. The above findings suggest that the searched bit combination based on Hessian sensitivity analysis is more effective and reasonable than uniform one, thereby boosting the performance of low-bit quantized models. Moreover, MSFT proves to be a powerful strategy for incrementally quanitzing and fine-tuning network in the context of mixed precision quantization.
@@ VII-D3 Binary Quantization
Finally, we analyze the binarized weight histograms of uniform, static and adaptive quantization schemes. The distributions of pre-trained weights, as shown in Fig. 9, display a notable disparity between the first and last convolutional layers. It is evident that uniform quantization gives the worst binarized results, failing to align with the real-valued weight distribution. Using a fixed integer set, static 1-bit quantization offers a closer match to the pre-trained weight distribution. However, it produces nearly identical binarized outcomes in the first and last layers ($0.07$ vs. $0.06$). For example, most weights in the final layer fall within the range $[-0.04,0.04]$, yet static quantization yields two binarized weights at ±$\pm 0.06$. In contrast, adaptive quantization can determine binary values for different layers according to their distributions, achieving the best performance. In the broader first layer, weights are mapped to ±$\pm 0.05$. Meanwhile, for the last layer with a more compact shape, it generates binarized weights of $-0.02$ and $0.02$. This illustrates the superiority of adaptive scheme over both the uniform and static method.
@@ VIII Conclusion
In this paper, we explore adaptive neural network quantization for lightweight speaker verification. Firstly, we introduce an innovative adaptive uniform precision quantization technique that leverages k-means clustering to dynamically generate quantization centroids specific to each network layer. To improve the performance of low-bit quantized models, a new algorithm that merges mixed precision quantization with a multi-stage fine-tuning (MSFT) strategy is further developed. Finally, we propose two distinct binary quantization schemes tailored for 1-bit scenario: the static and adaptive quantizers. Experimental results demonstrate that both ResNets and DF-ResNets effectively attain 4-bit uniform precision quantization with negligible performance loss. Furthermore, compared to the uniform precision approach, mixed precision quantization not only achieves better performance with a similar model size but also enables the generation of bit combination for any desirable model size. In addition, our newly designed 1-bit quantization schemes significantly enhance the performance of binarized models. The visualization of quantized weight distributions validates the superiority of our proposed quantization methods. Finally, an in-depth comparison with existing lightweight SV systems indicates that our resulting models substantially outperform earlier systems across various model size ranges.